OpenStudy (anonymous):

simplify 4^sqroot 80b^4c^3

OpenStudy (anonymous):

pass your test?

OpenStudy (anonymous):

\[\sqrt[4]{80b^4c^3}=\sqrt[4]{16\times b^4c^3}=\sqrt[4]{16}\times \sqrt[4]{b^4}\times \sqrt[4]{5c^3}\] \[=2b\sqrt[4]{5c^3}\]

OpenStudy (anonymous):

i got a 90

OpenStudy (anonymous):

not bad. but with all the help you should have got 100. we failed somewhere.

OpenStudy (anonymous):

its okay i just want to get this stuff done =/...im doing binomial and radical expressions now